The Magellanic Clouds Newsletter

An electronic exchange on Magellanic Clouds research

Edited by Eva K. Grebel and You-Hua Chu

          mcnews@astro.uiuc.edu
http://www.astro.uiuc.edu/mcnews/MCNews.html
  http://www.astro.uni-bonn.de/~mcnews/

No. 27 December 7, 1998

Contents

Proceedings announcement  1
Abstracts of 11 refereed papers  2
Abstract of 7 non-refereed papers 8
Thesis abstracts  12
Job opportunity  14

News

We are pleased to announce the availability of the proceedings of the following workshop:

Proceedings of the Bonn/Bochum-Graduiertenkolleg Workshop

The Magellanic Clouds and Other Dwarf Galaxies

held at the Physikzentrum Bad Honnef, Germany in January 1998

Tom Richtler & Jochen M. Braun (eds.)

Sternwarte der Universität Bonn, Auf dem Hügel 71, D–53121 Bonn, Federal Republic of Germany

This book addresses all those astrophysicists who take interest in the structure, kinematics and evolution of low-mass galaxies. It comprises nine reviews of experts in this field, plus a host of shorter contributions and posters, all reflecting the enormous progress that has been made in the recent past, partly due to superb new observing facilities (e.g. Keck, HST), but also to computing facilities that have made numerical modelling incredibly powerful to understand, e.g., stellar dynamics or hydrodynamics of the interstellar gas.

The book provides a quick overview of the recent achievements in dwarf galaxy research. This is naturally incentive for the understanding of stellar evolution and the interplay between the interstellar medium and star formation. Understanding the astrophysics of local (dwarf) galaxies is a necessary prerequisite for the comprehension of distant galaxies seen at earlier epochs.

The proceedings also reflect the harvest of some 5 years of continuous education and research in our graduate school, funded by the Deutsche Forschungsgemeinschaft and run by the Astronomical Institutes of the Universities of Bonn and Bochum. It is documented here how efficient and fruitful a coordinated educational programme can be that focusses on a fascinating, and therefore booming, field of research in contemporary astrophysics.

Printed version [79 contributions on 352 pages] Shaker Verlag, Aachen, ISBN 3-8265-4457-9

Full electronic version at URL http://www.astro.uni-bonn.de/~webgk/ws98/cover.html

For further information contact jbraun@astro.uni-bonn.de

Infos also available from the URL http://xxx.lanl.gov/abs/astro-ph/9811445

Abstracts of Refereed Papers

A Reanalysis of the Ultraviolet Extinction from Interstellar Dust in the Large Magellanic Cloud

K.A. Misselt, Geoffrey C. Clayton and Karl D. Gordon

Department of Physics and Astronomy, Louisiana State University, Baton Rouge, LA 70803-4001, USA

We have reanalyzed the Large Magellanic Cloud’s (LMC) ultraviolet (UV) extinction using data from the IUE final archive. Our new analysis takes advantage of the improved signal–to–noise of the IUE NEWSIPS reduction, the exclusion of stars with very low reddening, the careful selection of well matched comparison stars, and an analysis of the effects of Galactic foreground dust. Differences between the average extinction curves of the 30 Dor region and the rest of the LMC are reduced compared to previous studies. We find that there is a group of stars with very weak 2175 Å bumps that lie in or near the region occupied by the supergiant shell, LMC 2, on the southeast side of 30 Dor. The average extinction curves inside and outside LMC 2 show a very significant difference in 2175 Å bump strength, but their far–UV extinctions are similar. While it is unclear whether or not the extinction outside the LMC 2 region can be fit with the relation of Cardelli, Clayton & Mathis (CCM), sightlines near LMC 2 cannot be fit with CCM due to their weak 2175 Å bumps. While the extinction properties seen in the LMC lie within the range of properties seen in the Galaxy, the correlations of UV extinction properties with environment seen in the Galaxy do not appear to hold in the LMC.

Accepted by: The Astrophysical Journal

For preprints, contact misselt@physics.lsu.edu
Also available from the URL http://www.phys.lsu.edu/students/misselt/pubs.html

Supernova Remnants in the Magellanic Clouds II: SNR Breakouts from N11L and N86

Rosa Murphy Williams^{1}, You-Hua Chu^{1}, John R. Dickel^{1}, R. Chris Smith^{2}, Douglas K. Milne^{3} and P. Frank Winkler^{4}

1 Astronomy Department, University of Illinois, University of Illinois at Urbana-Champaign, Urbana, IL 61801, USA 2 Cerro Tololo Inter-American Observatory, La Serena, Chile 3 Australia Telescope National Facility, CSIRO, Box 76, Epping, New South Wales 2121, Australia 4 Department of Physics, Middlebury College, Middlebury, VT 05753, USA

The evolution of a supernova remnant (SNR) is heavily influenced by the interstellar conditions surrounding the remnant. This is particularly true in cases where the SNR is breaking out into a low-density area in the surrounding medium. We examine two promising candidates for the study of SNR breakouts in the Large Magellanic Cloud (LMC): N11L and N86. The remnant N11L has a filamentary extension which interrupts the shell; to the north of this extension, we find a region of diffuse radio and X-ray emission that shows only faint filaments in optical images. The discontinuous distribution of velocities in the shell material, and the apparent flattening of the radio spectral index in the outflow region, suggest substantial turbulence in the outflowing material and clumpiness in the surrounding ISM. N86, on the other hand, shows a large breakout to the north, as well as several smaller possible outflows around the shell. We find that the northern breakout has a well-defined spherical expansion pattern, faint diffuse X-ray emission, and a highly filamentary optical structure. Our velocity data indicate that material breaking out to the north is expanding at a maximum of 100 km s^{−1}. The consequences of these breakouts on the parent remnants are discussed: N11L appears to have a lower thermal energy, by an order of magnitude, than other LMC remnants used for comparison. N86, on the other hand, shows a thermal energy fairly similar to the comparison SNRs, perhaps due to a more gradual loss of hot gas. The implications of the breakout structures for the surrounding medium are also discussed. The breakout in N11L coincides with a possible low-density cavity which is enclosed in a shell structure on the western edge of the N11 H II complex. The less dense shell of N86, and the more distributed pattern of the breakouts, suggest a relatively low-density ISM with substantial local density variations.

Accepted by: The Astrophysical Journal (April 1 issue)

For preprints, contact rosanina@astro.uiuc.edu
Also available from the URL http://www.astro.uiuc.edu/~rosanina/n11n86.html

Structure and Evolution of Hot Gas in 30 Dor

Q. Daniel Wang^{1}

1 Northwestern University, 2145 Sheridan Road, Evanston, IL 60208-3112, USA

We have investigated the structure and evolution of hot gas in the 30 Dor nebula, based on recent X-ray observations. Our deep ROSAT HRI image shows that diffuse X-ray emission arises in blistershaped regions outlined by loops of HII gas. X-ray spectroscopic data from ASCA confirm the thermal nature of the emission and indicate that hot gas temperature decreases from the core to the halo of the nebula. The structure of the nebula can be understood as outflows of hot and HII gases from the parent giant molecular cloud of the central OB association. The dynamic mixing between the two gas phases is likely responsible for the mass loading to the hot gas, as required to explain the observed thermal structure and X-ray luminosity of the nebula. Such processes should also be important in the formation of similar giant HII regions and in their subsequent evolution into supergiant bubbles or galactic chimneys.

Accepted by: The Astrophysical Journal Letters (Jan. 10 issue) For preprints, contact wqd@nwu.edu

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9811031

Physical Structure of Small Wolf-Rayet Ring Nebulae

You-Hua Chu^{1}, Kerstin Weis^{2}, Donald R. Garnett^{3}

1 Astronomy Department, University of Illinois, 1002 W. Green Street, Urbana, IL 61801, USA 2 Institut für Theoretische Astrophysik, Tiergartenstr. 15, D-69121 Heidelberg, Germany 3 Steward Observatory, University of Arizona, Tucson, AZ 85721, USA

We have selected the seven most well-defined WR ring nebulae in the LMC (Br 2, Br 10, Br 13, Br 40a, Br 48, Br 52, and Br 100) to study their physical nature and evolutionary stages. New CCD imaging and echelle observations have been obtained for five of these nebulae; previous photographic imaging and echelle observations are available for the remaining two nebulae. Using the nebular dynamics and abundances, we find that the Br 13 nebula is a circumstellar bubble, and that the Br 2 nebula may represent a circumstellar bubble merging with a fossil main-sequence interstellar bubble. The nebulae around Br 10, Br 52, and Br 100 all show influence of the ambient interstellar medium. Their regular expansion patterns suggest that they still contain significant amounts of circumstellar material. Their nebular abundances would be extremely interesting, as their central stars are WC5 and WN3-4 stars whose nebular abundances have not been derived previously. Intriguing and tantalizing implications are obtained from comparisons of the LMC WR ring nebulae with ring nebulae around Galactic WR stars, Galactic LBVs, LMC LBVs, and LMC BSGs; however, these implications may be limited by small-number statistics. A SNR candidate close to Br 2 is diagnosed by its large expansion velocity and nonthermal radio emission. There is no indication that Br 2’s ring nebula interacts dynamically with this SNR candidate.

Accepted by: The Astronomical Journal (March issue) For preprints, contact chu@astro.uiuc.edu

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9812076 or
      http://www.astro.uiuc.edu/~chu/publications/preprints/LMC WR.ps

A Model For the Outer Rings of SN 1987A

Noam Soker

Dept. of Physics, Univ. of Haifa, Israel

I propose a model for the formation of the two outer rings of SN 1987A. The main new ingredient is a short-lived, few×10 years, intermediate wind, of velocity ∼ 100 km s^{−1} and total mass of ∼ 10^{−3}M_{⊙}−few×10^{−2}M_{⊙}, which is concentrated near the equatorial plane. This intermediate wind was formed during the few orbital periods when a binary companion entered the envelope of the progenitor of SN 1987a, when the latter was a red supergiant. The intermediate wind formed a hollowed torus with a horseshoe-shaped cross section. The inner regions at the ends of the horseshoe torus are the two outer rings of SN 1987A. The other regions of the torus were cleaned by the fast wind blown after the progenitor turns into a blue giant. The proposed scenario accounts for: (i) slow motion of the outer rings; (ii) the high density contrast of the outer rings to their surroundings; (iii) the presence of only two outer rings; (iv) the displacement of the rings relative to the central star; (v) the fact that no planetary nebulae have been observed to have a single pair of outer rings.

Accepted by: MNRAS

For preprints, contact soker@physics.technion.ac.il
Also available from http://xxx.lanl.gov/abs/astro-ph/9811151

Models of X-ray Photoionization in LMC X-4: Slices of a Stellar Wind

B. Boroson^{1}, T. Kallman^{1}, R. McCray^{2}, S.D. Vrtilek^{3}, and J. Raymond^{3}

1 Goddard Space Flight Center, Greenbelt, MD 20771, USA 2 JILA, Campus Box 440, University of Colorado, Boulder, CO 80309, USA 3 Center for Astrophysics, 60 Garden Street, Cambridge, MA 02138, USA

We show that the orbital variation in the UV P Cygni lines of the X-ray binary LMC X-4 results when X-rays photoionize nearly the entire region outside of the X-ray shadow of the normal star. We fit models to Goddard High Resolution Spectrograph (GHRS) observations of N V and C IV P Cygni line profiles. Analytic methods assuming a spherically symmetric wind show that the wind velocity law is well-fit by v≈(1–1/r)^{β}, where β is likely 1.4–1.6 and definitely < 2.5. Escape probability models can fit the observed P Cygni profiles, and provide measurements of the stellar wind parameters. The fits determine L_{x}/Ṁ=2.6 ± 0.1 × 10^{43} erg s^{−1} M_{⊙}^{−1} yr, where L_{x} is the X-ray luminosity and Ṁ is the mass-loss rate of the star. Allowing an inhomogeneous wind improves the fits. IUE spectra show greater P Cygni absorption during the second half of the orbit than during the first. We discuss possible causes of this effect.

Accepted by: The Astrophysical Journal

For preprints, contact bboroson@falafel.gsfc.nasa.gov
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9811267

RXT E Observations of LMC X-1

P.C. Schmidtke^{1}, A.L. Ponder^{1}, and A.P. Cowley^{1}

1 Department of Physics & Astronomy, Arizona State University, Tempe, AZ, 85287-1504, USA

The luminous X-ray binary and black-hole candidate LMC X-1 has been observed with the Rossi X-ray Timing Explorer (RXT E) to search for quasi-periodic oscillations (QPO), previously reported in its high state. The source was observed monthly in a series of nine observations. Analysis of the temporal variations shows no evidence for QPO or other periodic changes, but correlations between the high-energy light curve and hardness of the spectrum are described. Spectral fits with two-component models demonstrate that the hardness variations come from changes in the intensity of the high-energy power-law tail.

Accepted by: The Astronomical Journal

For preprints, contact Paul.Schmidtke@asu.edu

The Dependence of the Soft X-ray Properties of LMXBs on the Metallicity of Their Environment

J.A. Irwin^{1} and J.N. Bregman^{1}

1 University of Michigan, Department of Astronomy, 909 Dennison Building, Ann Arbor, MI 48109-1090, USA

We determine the X-ray spectral properties of a sample of low-mass X-ray binaries (LMXBs) which reside in globular clusters of M31, as well as five LMXBs in Galactic globular clusters and in the Large Magellanic Cloud using the ROSAT PSPC. We find a trend in the X-ray properties of the LMXBs as a function of globular cluster metallicity. The spectra of LMXBs become progressively softer as the metallicity of its environment increases. The one M31 globular cluster LMXB in our sample which has a metallicity greater than solar has spectral properties similar to those of LMXBs in the bulge of M31, but markedly different from those which reside in low metallicity globular clusters, both in M31 and the Galaxy. The spectral properties of this high metallicity LMXB is also similar to those of X-ray faint early-type galaxies. This lends support to the claim that a majority of the X-ray emission from these X-ray faint early-type galaxies results from LMXBs and not hot gas, as is the case in their X-ray bright counterparts.

Accepted by: Astrophysical Journal Letters

For preprints, contact jirwin@astro.lsa.umich.edu
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9811063

Magellanic Cloud X-ray Sources: III. Completion of a ROSAT Survey

P.C. Schmidtke^{1}, A.P. Cowley^{1}, J.D. Crane^{1}, V.A. Taylor^{1}, T.K. McGrath^{1}, J.B. Hutchings^{2}, and David Crampton^{2}

1 Department of Physics & Astronomy, Arizona State University, Tempe, AZ, 85287-1504, USA 2 Dominion Astrophysical Observatory, National Research Council of Canada, Victoria, B.C., V8X 4M6, Canada

This paper concludes a series of three papers presenting ROSAT High-Resolution Imager (HRI) observations of unidentified Einstein and serendipitous ROSAT X-ray sources in the direction of the Magellanic Clouds. Accurate positions and fluxes have been measured for these sources. Optical photometry and spectroscopy were obtained to search for identifications in order to determine the physical nature of these sources. The present paper includes new data for 24 objects; identifications are given or confirmed for 30 sources. For six sources optical finding charts showing the X-ray positions are provided. The results from this program are summarized, showing the populations of luminous X-ray sources in the Magellanic Clouds are quite different from those in the Galaxy.

Accepted by: The Astronomical Journal

For preprints, contact Paul.Schmidtke@asu.edu

The MACHO Project Large Magellanic Cloud Variable Star Inventory. VIII. The Recent Star Formation History of the LMC

from the Cepheid Period Distribution

C. Alcock^{1}, R.A. Allsman^{2}, D. Alves^{1}, T.S. Axelrod^{2}, A.C. Becker^{3}, D.P. Bennett^{1}, D.F. Bersier^{2}, K.H. Cook^{1}, K.C. Freeman^{2}, K. Griest^{4}, M.J. Lehner^{4}, S.L. Marshall^{1},

D. Minniti^{1},^{5}, B.A. Peterson^{2}, M.R. Pratt^{3}, P.J. Quinn^{6}, A.W. Rodgers^{2}, C.W. Stubbs^{3}, W. Sutherland^{7}, A. Tomaney^{3}, T. Vandehei^{4}, D.L. Welch^{8}

1 Lawrence Livermore National Laboratory, Livermore, CA 94550, USA 2 Mt Stromlo Observatory, Weston, ACT 2611, Australia 3 Department of Astronomy, University of Washington, Seattle, WA 98195, USA 4 Department of Physics, University of California, San Diego, CA 92093, USA 5 Departmento de Astronomia y Astrofisica, P. Universidad Catolica, Casilla 104, Santiago 22, Chile 6 European Southern Observatory, D-85748 Garching bei München, Germany 7 Department of physics, University of Oxford, Oxford OX1 3RH, U.K. 8 Department of Physics and Astronomy, McMaster University, Hamilton, Ontario L8S 4M1, Canada

We present an analysis of the period distribution of ∼ 1800 Cepheids in the Large Magellanic Cloud, based on data obtained by the MACHO microlensing experiment and on a previous catalogue by Payne-Gaposchkin. Using stellar evolution and pulsation models, we construct theoretical periodfrequency distributions that are compared to the observations. These models reveal that a significant burst of star formation has occurred recently in the LMC (∼ 1.15×10^{8} years). We also show that during the last ∼ 10^{8} years, the main center of star formation has been propagating from SE to NW along the bar. We find that the evolutionary masses of Cepheids are still smaller than pulsation masses by ∼ 7 % and that the red edge of the Cepheid instability strip could be slightly bluer than indicated by theory. There are ∼ 600 Cepheids with periods below ∼ 2.5 days cannot be explained by evolution theory. We suggest that they are anomalous Cepheids; a number of these stars are double-mode Cepheids.

Accepted by: The Astronomical Journal

For preprints, contact bersier@mso.anu.edu.au
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9811240

Microlensing in the Small Magellanic Cloud: lessons from an N-body simulation

D.Graff^{1}, L.Gardiner^{2}

1 The Ohio State University, Departments of Astronomy and Physics, Columbus, OH 43210, USA 2 Sun Moon University, Dept. of International Education, Tangjeongmyeon, Asankun, Chungnam, Republic of Korea

We analyse an N-body simulation of the Small Magellanic Cloud (SMC), that of Gardiner & Noguchi (1996) to determine its microlensing statistics. We find that the optical depth due to selflensing in the simulation is low, 0.4 × 10^{−7}, but still consistent (at the 90 % level) with that observed by the EROS and MACHO collaborations. This low optical depth is due to the relatively small line of sight thickness of the SMC produced in the simulation. The proper motions and time scales of the simulation are consistent with those observed assuming a standard mass function for stars in the SMC. The time scale distribution from the standard mass function generates a significant fraction of short time scale events: future self-lensing events towards the SMC may have the same time scales as events observed towards the Large Magellanic CLoud (LMC). Although some debris was stripped from the SMC during its collision with the LMC about 2 × 10^{8} yr ago, the optical depth of the LMC due to this debris is low, a few ×10^{−9}, and thus cannot explain the measured optical depth towards the LMC.

Submitted to: MNRAS For preprints, contact graff.25@osu.edu

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9811394

Abstracts of Non-Refereed Papers

Molecular Gas in the Magellanic Clouds

Mónica Rubio ^{1}

1 Departmento de Astronomía, Universidad de Chile, Casilla 36-D, Santiago, CHILE

The molecular gas content in the Magellanic Clouds has been studied, with different spatial coverage and resolution, through obervations of CO(1-0) line emission. In LMC and SMC the molecular gas is dominated by clouds whose properties are different to those of their Galactic counterparts. The relation between the intensity of CO emission and molecular hydrogen column density, or conversion factor X, is different than that of molecular clouds in our Galaxy and depends on the ambient physical conditions. Studying the molecular gas through observations in the H_{2} emission line may prove an alternative way to determine the molecular content associated to star forming regions in the Magellanic Clouds. In particular, results obtained towards 30 Doradus in the LMC are presented.

Invited talk to appear in: IAU Symp. 190, New Views of the Magellanic Clouds, Eds. Chu et al., ASP Conference Series

For preprints, contact mrubio@das.uchile.cl

The Metallicity Effect on the Cepheid P-L relation from SMC Cepheids

Jesper Storm^{1}, Bruce W. Carney^{2}, and Anne M. Fry^{3}

1 Astrophysikalisches Institut Potsdam, Germany 2 University of North Carolina at Chapel Hill, NC, USA 3 Case Western Reserve University, OH, USA

We present results based on a Baade-Wesselink analysis of Cepheids in the Small Magellanic Cloud. The Baade-Wesselink analysis provides individual luminosities for these metal-poor Cepheids which combined with recent Baade-Wesselink results from Gieren et al. (1998) on solar metallicity Galactic Cepheids constrain the metallicity effect on the zero-point of the Cepheid P-L relation. A preliminary analysis leads to an effect of ΔM_{V} /Δ[Fe/H] = −0.45 ± 0.15, metal-rich Cepheids being brighter, in good agreement with several recent independent determinations. An effect of this magnitude reduces significantly the current disagreement between the long and the short distance estimates to the LMC, and favors a shorter value.

To appear in: ”Harmonizing Cosmic Distance Scales in a Post-Hipparcos Era”, D. Egret and A. Heck Eds., ASP Conf. Series, 1999 For preprints, contact jstorm@aip.de

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9811376

First results from the EROS-II microlensing experiment

O. Perdereau^{1} (on behalf of the EROS-2 collaboration)

1 LAL, IN2P3-CNRS and Univ. de Paris-Sud BP 34 F-91898 Orsay Cedex, France

EROS-II is a second generation microlensing experiment. The experimental setup, in operation at the European Southern Observatory (ESO) at La Silla (Chile) since mid-1996 is briefly described together with its scientific objectives. The first results from our microlensing searches towards the Small Magellanic Cloud (SMC) and towards the Galactic Plane are presented. We also give some results from a dedicated campaign which took place in 96-97, and aimed at studying Magellanic Cepheids systematically. We conclude with an overview of the semi-automated supernovæ search.

Proceedings of the IIIrd Stromlo Symposium, to be published by ASP, B. Gibson, T. Axelrod and M. Putman (eds)

For preprints, contact perderos@lal.in2p3.fr
Also available from the URL http://www.lal.in2p3.fr/EROS
or by anonymous ftp at ftp.lal.in2p3.fr/Eros/Papiers/perdereau.ps.gz

Near-IR Catalogue of the Magellanic Clouds: DENIS (Deep Near-IR Southern Sky Survey

M.R. Cioni^{1}, H.J. Habing^{1}, C. Loup^{2}, M.A.T. Groenewegen^{3}, N. Epchtein^{4} and the DENIS Consortium^{5}

1 Sterrewacht Leiden, Postbus 9513, 2300 RA Leiden, Nederland 2 IAP, 98 bis Boulevard Arago, F-75014 Paris, France 3 MPI für Astrophysik, Karl-Schwarzschild-Straße 1, D-85748 Garching, Germany 4 Obs. de Côte d’Azur, 06304 Nice Cedex 04, France 5 See aknowledgments

We report here on a preliminary analysis of the DENIS new IR point source catalogue of the Magellanic Clouds. After subtracting the contribution of foreground stars we interpret the colour-colour diagrams and colour-magnitude diagrams that characterize the Asymptotic Giant Branch (AGB). The AGB stars cause a clear bump in the luminosity function and they distribute in two distinct regions according to the spectral type (C-rich and O-rich) in the colour-luminosity diagram.

To appear in: IAU Symposium 192, “The Stellar Content of Local Group Galaxies” (Cape Town 7-11 September 1998)

For preprints, contact mrcioni@strw.leidenuniv.nl

Effect of the Magellanic Clouds on the Milky Way disk and VICE VERSA

Martin D. Weinberg^{1}

1 Department of Physics & Astronomy, University of Massachusetts, Amherst, MA 01003-4525, USA

The satellite-disk interaction provides limits on halo properties in two ways: (1) physical arguments motivate the excitation of observable Galactic disk structure in the presence of a massive halo, although precise limits on halo parameters are scenario-dependent; (2) conversely, the Milky Way as a whole has significant dynamical effect on LMC structure and this interaction also leads to halo limits. Together, these scenarios give strong corroboration of our current gravitational mass estimates and suggests a rapidly evolving LMC.

To appear in: the Third Stromlo Symposium: The Galactic Halo (ASP Conference Series), in press.

For preprints, contact weinberg@phast.umass.edu
Also available from the URL http://www-astro.phast.umass.edu/~weinberg/stromlo
or see astro-ph/9811204

The H I fine structure of HVC 187 near NGC 3783: gas in the leading bridge of the Magellanic System

B.P. Wakker^{1}, B.D. Savage^{1}, T.A. Oosterloo^{2}, M.E. Putman^{3}

1 Department of Astronomy, University of Wisconsin, Madison, WI 53706, USA 2 CNR, Istituto di Fisica Cosmica, Milan, Italy 3 Mount Stromlo and Siding Springs Observatories, Australia

We present an analysis of high-resolution H I data of one of the cores of HVC 187 (Wakker & van Woerden 1991), HVC287+22+240. Structure is present down to the lowest-measurable scale (∼1 arcmin) and several concentrations appear to be unresolved. Most of the cores seen at low resolution break up into smaller subcores at higher resolution. The typical volume density and pressure are estimated to be ∼30 R^{−1} D_{kpc}^{−1} cm^{−3} and ∼18000 R^{−1} D_{kpc}^{−1} K cm^{−3}, respectively, where R is the resolution in arcmin and D_{kpc} the unknown distance in kpc.

To appear in: Proceedings of the Third Stromlo Symposium, Canberra August 1998, eds. B.K. Gibson, T.S. Axelrod & M.E. Putman (ASP Conf. Ser.)

Observational evidence for the accretion of low-metallicity gas onto

the Milky Way: metallicity, physical conditions and distance limit for HVC complex C

B.P. Wakker^{1}, J.C. Howk^{1}, B.D. Savage^{1}, S.L. Tufte^{1}, R.J. Reynolds^{1} H. van Woerden^{2}, U.J. Schwarz^{2}, R.F. Peletier^{3}

1 Department of Astronomy, University of Wisconsin, Madison, WI 53706, USA 2 Rijks Universiteit Groningen, The Netherlands 3 University of Durham, UK

We present observations of the (field of the) Seyfert galaxy Mark 290, which probes the highvelocity cloud (HVC) complex C, one of the largest HVCs (Wakker & van Woerden 1991). We find that this object has a metallicity of 0.094±0.020_{−0.019}^{+0.022} times solar. We determine a lower limit to its distance of 5 kpc (z>3.5 kpc). If the gas is in thermal equilibrium with a hot halo, then it is also likely that D<30 kpc, putting the HVC in the Galactic Halo, as was the case for HVC complex A (for which z=2.5–7 kpc; van Woerden et al. 1998).

We find that, on this sightline, H^{+} represents 23±10% of the hydrogen in the HVC. The total gaseous mass is 2.0×10^{6} (D/5 kpc)^{2} M_{⊙} and, depending on whether the space velocity is completely radial or vertical, the HVC represents a downward mass flux of ∼0.036–0.083 (D/5 kpc) M_{⊙} yr^{−1}, or ∼2.9–6.7×10^{−3} (D/5 kpc)^{−1} M_{⊙} yr^{−1} kpc^{−2}.

The low metallicity and large mass suggest that complex C is unlikely to be part of a Galactic Fountain, but rather represents accreting low-metallicity material. It may be a present-day analogue of the damped Lyα absorbers seen in QSO spectra. Our abundance result provides the first direct observational evidence for the infall of low-metallicity gas on the Milky Way, required in models of galactic chemical evolution.

It remains to be seen whether ultimately complex C is a remnant of the formation of the Milky Way (Oort 1970), a gas cloud orbiting the Galaxy (Kerr & Sullivan 1969), a Local Group object (Verschuur 1969, Blitz et al. 1996), the result of tidal interactions between the Magellanic System and the Galaxy >3 Gyr ago (an “Old Magellanic Stream”; Toomre, quoted in Kerr & Sullivan 1969), or was formed when hot, ionized intergalactic gas was compressed by the motion of the Milky Way (Silk et al. 1987).

To appear in: Proceedings of the High-Velocity Cloud workshop, Canberra August 1998, eds. B.K. Gibson & M.E. Putman (ASP Conf. Ser.)

Thesis Abstracts

Studies of Interstellar Molecular Gas in the Magellanic Clouds

Arto Heikkilä

Onsala Space Observatory, SE-439 92 Onsala, Sweden Present address: Observatory, P.O. Box 14, FIN-000 14 University of Helsinki, Finland

Results of observational studies of interstellar molecular gas clouds in the Large Magellanic Cloud (LMC) and the Small Magellanic Cloud (SMC), two nearby metal-poor and dust-poor dwarf galaxies, are presented. Spectral line radiation emitted by interstellar molecules has been used to probe the physical state and the chemical composition of individual molecular clouds. In addition, observations of the C^{18}O/C^{17}O ratio in a few Galactic and extragalactic sources are presented.

Two molecular species, deuterated formyl ion (DCO^{+}) and hydrogen sulphide (H_{2}S), were detected for the first time in an extragalactic source (N159W in the LMC). Moreover, a definite detection of C^{17}O outside the Local Group has been obtained.

The density and the temperature of the molecular gas in the Magellanic Clouds are estimated to be typically 10^{4}–10^{6} cm^{−3} and 15–30 K, respectively. The abundances of trace molecules (relative to H_{2}) in clouds in the LMC are typically five to twenty times lower than in Galactic disc clouds. These underabundances are most likely caused by a combination of the lower metallicity and a more rapid photodissociation in the LMC. The effects of the metallicity and far-ultraviolet radiation are further emphasized in two of the clouds, where trace molecules are even less abundant: the cloud (N27) studied in the SMC and a cloud located near the centre of the 30 Doradus nebula in the LMC.

Several features, most notably the abundance of C_{2}H as well as the correlations of the HCO^{+}/CO and the C_{2}H/CO ratios with the star-formation activity, suggest that the chemistry in molecular clouds in the LMC and the SMC largely resembles that in Galactic photon-dominated regions.

The gas-phase ^{12}C/^{13}C and ^{32}S/^{34}S abundance ratios are found to be similar to those measured in Galactic clouds. In particular, no trend of a higher ^{12}C/^{13}C ratio due to the low metallicity is discernible. On the other hand, the ^{18}O/^{17}O abundance ratio, the average in the LMC is estimated to be 1.6±0.3, is significantly lower than in Galactic clouds (by a factor of two) and centres of starburst galaxies (by a factor of five). Provided that the current understanding of stellar nucleosynthesis applies, a low ^{18}O/^{17}O abundance ratio indicates that massive stars have contributed only little to the enrichment of the interstellar medium. This suggests that the star-formation history in the LMC has been characterised by a steep initial mass function together with a low average star-formation rate.

PhD thesis completed at Chalmers University of Technology, Göteborg, Sweden in October 1998. Supervisor: Dr. Lars E.B. Johansson

   For copies, contact: arto@obs13.astro.helsinki.fi

On the Binary Frequency of a Complete Sample of Magellanic, WC

Wolf-Rayet Stars and a Spectroscopic Study of WC Binary Colliding Winds

P. Bartzakos

Département de physique, Université de Montréal, P.O. Box 6128, Station “Centre-Ville”, Montreal, P.Q., H3C 3J7, Canada

A spectroscopic and photometric study of all 24 known WC/WO Wolf-Rayet stars of the Magellanic Clouds is made in an effort to determine the binary frequency of WR-stars in environments of low metallicity. This is followed by an investigation of colliding winds in those stars revealed to be binary.

Low metallicity has been postulated in the literature as the prime cause for the low WR/O number ratio in the Clouds. This follows from the fact that progenitor O-star winds at low Z have insufficient strength to repel the outer hydrogen envelope of all but the most massive, luminous O-stars. This removal is needed to produce a WR-star. Low-to-medium mass O-stars in low-Z environments require a companion to reduce their gravitational potential and thus ease the removal of the outer layers. By studying absolute magnitudes, line equivalent widths, and radial velocity variability, a binary frequency of only thirteen percent is found for the Large Magellanic Cloud. (The Small Cloud contains only one WC-star.) This is far less than the 52%±14% frequency predicted by theory. The stellar winds appear powerful enough to accomplish the task themselves, even in a low-metallicity milieu.

As O-star winds are also strong, they should collide with the WR winds in WR+O binaries to form a bow-shaped shock front. The shocked material is heated to high temperature near its head, after which it flows along a shock cone that wraps around the weaker-wind O-star. Rapid cooling of the dense shock zone yields radiation much like that seen from the WR wind itself. Such excess emission is seen in many WR emission lines; it generally produces double-peaked spectral profiles that can be described by the model of Lührs. These profiles depend on orbital phase, orbital plane inclination, streaming velocity of the shocked material and cone opening angle. Such profiles are modelled for Ciii 5696 Å in Br22 (the only WC/WO system in the MC’s to reveal significant Ciii excess) in an attempt to determine these parameters. However, a Lührs’ model fit to the observed Ciii excess in Br22 gives only qualitative agreement. A new global approach proves more successful. The excess profile is characterised by the full width at half-maximum and mean velocity. A fit of the simplified Lührs’ model in this case is used in an attempt to determine the cone parameters and orbital inclination. However, because of observational errors, the solution appears to be numerically degenerate in orbital inclination. The method, however, shows promise if more high-quality data can be used to reduce the noise. The presence of excess emission in the form of Lührs’ profiles in the Civ 5808 Å line, present in all binaries of the sub-sample, shows that the colliding wind phenomenon is universal amongst close WC/WO binary stars.

Thesis accepted on November 13, 1998, at the Université de Montréal, under the direction of Professor A.F.J. Moffat.

Job Opportunity

Postdoctoral Fellowship at the University of Bonn

The Graduiertenkolleg ”The Magellanic Clouds and Other Dwarf Galaxies”, jointly run at the Universities of Bonn and Bochum, funded by the Deutsche Forschungsgemeinschaft, invites applications for a postdoctoral fellowship. The Fellow is meant to work at Bonn University.

The successful applicant will find her/himself in a pleasant scientific environment, closely collaborating with graduate students on related topics. She/he should be interested in galaxy research, in particular in the investigation of dwarf galaxies. Members of the Graduiertenkolleg currently pursue both theoretical and observational studies in the field of galaxy reseach. They have regular access to major radio and optical telescopes around the world, and also utilize satellite observatories (e.g. in X-ray astronomy).

The major goal of this graduate or research school is to furnish comprehensive scientific experience and education of graduate students in astrophysics, while working in a team. The scientific focus is on the star formation history and evolution, the interstellar medium, and the structure and kinematics of the Magellanic Clouds and of more distant dwarf galaxies.

For further information, please, see also our grad school home page:

                         http://www.astro.uni-bonn.de/ webgk

The DFG will continue funding for the next 3 years, and it is foreseen that the fellowship will be available as of 1st of January next year. The appointment is to the end of 2000. The monthly salary is DM 2490.- and 2690.- (depending on age), plus DM 200.- for ex- penditures, plus a DM 400.- allowance for married persons.

Applications should be sent to the Spokesman of the Graduiertenkolleg:

Prof. Dr. Uli Klein

Radioastronomisches Institut der Universitaet Bonn

Auf dem Huegel 71

D-53121 Bonn, Germany.